Report of Activities on the RVIB N.B. Palmer
25 August 2001

The marginal ice edge zone, the location where there is a transition from pack ice to open water, in the winter is ill defined. In the Antarctic Western Peninsula Region north of our survey area, it can be spread out over 10's of miles under one wind condition (westerlies or south westerlies) or more compacted during another (north easterlies). During the morning of 26 August, we steamed to the northeast from broad-scale station 4 through thinner ice with more open leads headed for the marginal ice zone, where a survey of birds and marine mammals took place later in the day. We are now steaming towards the Bismark straits and later in the night we will enter the Gerlache Straits. Our destination is Paradise Harbor, which will be the location where the acoustic calibrations will take place. Our current position at 2021 is -65° 01.184S; -65° 25.58W and it is snowing. The wind is out of the north at 23 to 25 kts, the temperature is -1.6°C, and the barometer is holding steady at 987.7 mlb.

Shortly after midnight on 25 August, the Palmer arrived on station #15 and BIOMAPER-II was brought on board to allow the CTD cast to be done. Ice and wind conditions were such that it would have been difficult to have "parked" the towed body at its normal on-station depth (around 30 m) and to position the ship to do the CTD. It was again deployed once the CTD was back on board and the towyo to station #16 was easier than to station #15 because the pack ice became thinner and leads more frequent. At station #16, BIOMAPER-II was recovered. The station work consisted of a CTD, 10-m2 MOCNESS tow, ROV under-ice survey, ice collection, and a Plummet Net cast. The 10-m2 MOCNESS again had its problems with the ice. At the time of the start of the tow in a fair-sized lead, visibility was good in spite of low clouds. But during the tow, a heavy snowfall started and the wind picked up, significantly reducing the visibility. The tow was completed about 0950, but the ending was bad. The ship ran out of the lead before the last net had finished fishing and was closed. As a result the last net caught some ice and that made recovery of the net system, a real bear. In addition, the nets again froze the minute they got above the water and become stiff as a board. This time a rig was set up to use the tugger to help pull the nets up, but it was very time consuming and the working conditions with the blowing snow in the 30 kt winds and temperatures around -15° C made it very difficult net recovery. A 1-m2 MOCNESS tow was scheduled, but was cancelled because the lead that the tow was going to be done in had closed significantly and was too short. In addition, calibration of the acoustic systems was cancelled because of the weather conditions. The ROV survey under the ice was possible and produced a lot of excitement when the VPR cameras showed that present were some of the largest larval krill concentrations yet seen associated with the pack ice. This generated the need to do an ice collection and to try and sample the krill larvae using a pumping system hose pushed through a coring hole to sample the water under the ice. Work at the station ceased about 1830 and the long steam to the ice edge zone via station #4 began. As a result of the day-long series of measurements made at station #16, no bird or marine mammal surveys were done.

John Klinck reports that on August 25, the CTD group did 2 CTD casts and one XBT. The CTD stations were on the middle part of the 420 line, which is being re-sampled. The XBT was done between stations #15 and 16.

The XBT had multiple temperature maxima of 1.7 between 250 and 320 m indicating the presence of oceanic water in the region.

Station #15 (541 m) has a mixed layer to 75 m with little gradient. The temperature maximum is about 1.75°C at 240 m. The O2 minimum is at 200 m. Temperature reversals can be seen from the lower pycnocline to the bottom.

Station #16 (521 m) has a mixed layer to about 85 m with no gradients. A gentle pycnocline extends from 100 to 250 m with the temperature maximum at about 280 m with a value of 1.77°C. Numerous temperature reversals occurred between 250 m and the bottom. The O2 minimum is at 250 m.

More interesting is the change in water properties between the two samples, the first of which were on yearday 212, about 24 days ago. Originally, very warm deep water (1.88°C) was found only at station #16. Now both #15 and 16 show oceanic water with deep maximum temperatures of 1.77 and 1.74, respectively. The original water may have diffused/mixed or moved laterally or it may have lost its heat vertically into the cold mixed layer. The ice was much thinner at this place compared to offshore, which may be related or it may be coincidence. In addition, there is a cooler, higher oxygen water mass on the shelf at the bottom. This water comes from 700 to 800 m depth off the shelf break, based on the T, S, and O2 properties; it looks like Lower CDW. The path and mechanism by which this water rises 300 m and moves onto the shelf remains to be resolved.

Kendra Daly reports that on back on 22 August, the Daly/Torres group did a Tucker trawl to about 500 m at station #71 at the shelf break and collected 4 myctophid fishes, a midwater shrimp, Pasiphaea, several species of copepods, including female Paraeuchaeta carrying an egg sac and a large red copepod that may be Megacalanus sp., ostracods, chaetognaths, a siphonophore, an Atolla medusae, and several larval Euphausia frigida. During a revisit to the third transect line on 23 August, a Plummet net at station #14 collected many larval E. superba in the upper 100 m, where the 200 kHz acoustic trace indicated there was significant backscatter with intermittent stronger aggregations. A second net tow between 300 and 425 m collected Thysanoessa macrura and several species of copepods.

A final Plummet net tow on August 25 at station #16 collected a few E. superba furcilia, naked and shelled pteropods, a few Thysanoessa, and a ctenophore in the upper 100 m. A 200 kHz layer detected between 20 and 60 m may have been due in part to the hard-shelled pteropods, which give a strong acoustic signal. Although few furcilia were collected in the water column, large numbers were observed on the undersurface of sea ice by the ROV. In addition, large particles appeared to be sloughing off the undersurface of the sea ice. After ice cores were collected at this site, a pump system was used through the coring hole to collect water and furcilia at the ice-water interface. The water will be analyzed for size fractioned chlorophyll and particulate organic carbon and nitrogen and for autotrophic, heterotrophic and detrital components in collaboration with Scott Gallager to determine the food quality and quantity available to larval krill.

The Daly group also completed a larval and adult krill feeding experiment, a growth and molting rate experiment, and an experiment to measure the rate at which krill assimilate carbon and nitrogen from their ingested food.

Ana Sirovic report that on 24 August, she deployed 2 difar sonobuoys as the Palmer steamed to station #13. The first sonobuoy failed after only 4 minutes of transmission. The second one was monitored for 2 hours 2 minutes. Potential biological noise was heard on this buoy, but the recording needs to be reanalyzed before final decision on the origin of noise is made. No sonobouys were deployed on 25 August, since we were occupying station #16 for most part of the day and deployments would have resulted mainly in very loud ship noise.

BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
With the completion of the towyo section along transect 3, data collection with BIOMAPER-II came to an end for this cruise. All that remains to be done is the calibration of the transducers in a protected area before setting sail across the Drake Passage.

There was no 1-m2 MOCNESS tow on 25 August.

Cheers, Peter